OpenStudy (anonymous):

derivative of y=-cos^-1(1/x^3)

OpenStudy (anonymous):

d/dx(-cos^(-1)(1/x^3)) = -3/(sqrt(1-1/x^6) x^4)

OpenStudy (anonymous):

\[y=-\cos^{-1}\left(\frac{1}{x^3}\right)\\ \cos (-y)=\frac{1}{x^3}\\ \cos y=\frac{1}{x^3}\] Implicit differentiation yields \[-\sin y~\frac{dy}{dx}=-\frac{3}{x^4}\\ \frac{dy}{dx}=\frac{3}{x^4}\csc y\]